OpenStudy (anonymous):

If a = 3 and b = 5, find a/b=___?

OpenStudy (anonymous):

0.6

OpenStudy (anonymous):

these are the answer choices 2/5 3/5 5/3

OpenStudy (turingtest):

put them in your calculator and see which one is equal to 0.6

OpenStudy (nenadmatematika):

3/5

OpenStudy (turingtest):

3/5=6/10=0.6

OpenStudy (anonymous):

3/5

OpenStudy (turingtest):

more importantly, note how the answer is just found by putting the numbers for a and b in the proper places